# Approximating the set of Nash equilibria for convex games

### v2

@@ 5.2 Pollution control game
Let us consider the $N$-player pollution control game from environmental economics considered in [20]. There, each player represents a country that is, on one hand, seeking to maximing the net revenue from production of goods and services and, on the other hand, minimizing the environmental damage due to pollution. Pollution is assumed to be a proportional by-product of production and therefore the net revenues (gross revenue minus production cost) of each player, e.g., it can be expressed as a function of emissions. The strategy $x_{i}$ of country $i$, which is chosen from the space $\mathcal{X}_{i}=\mathbb{R}$, can be seen as the emissions of country $i$. Each country has its own revenue function $g_{i}:\mathcal{X}_{i}\to\mathbb{R}_{+}$ which is nonnegative and concave and depends only on the emissions of country $i$. However, the environmental damage due to emissions depends on the sum of all combined emissions which is stated by a convex damage cost function $d$. Each country tries to minimize its cost function $f_{i}$ (i.e. to maximize her total welfare given as the difference between the net revenue and the damage cost), which is given by $f_{i}(x):=d(x_{1}+...+x_{N})-g_{i}(x_{i})$. Additionally, two types of constraints can be considered. Firstly, each country $i$ can face an environmental constraint $x_{i}\in[0,T_{i}]$ where $T_{i}\geq 0$ is an exogenously given upper bound on emissions, e.g. from an international treaty. Secondly, countries can agree to jointly fix an upper bound for emissions, which is stated by a polyhedral constraint of the form $\sum_{i=1}^{N}\alpha_{i}x_{i}\leq T$, where $\alpha_{i}\in\mathbb{R}_{+}$ for $i=1,...,N,$ are positive coefficients and $T>0$ is an upper bound. In [20], the assumption of a unique Nash equilibrium is made and only those pollution control games are considered in their study. In contrast, with help of Theorem 3.1 (see also [5, Th. 2.6]) we can characterize the set of Nash equilibria as the set of all Pareto solutions of a certain vector optimization problem, regardless of uniqueness or finiteness of the set of equilibria. Furthermore, with Algorithm 4.7, we are able to numerically approximate the set of Nash equilibria. In the following, we present examples for games with $2$ or $3$ players, and where the set of Nash equilibria is nontrivial.

### v4

@@ 5.3 Pollution control game
Let us consider the $N$-player pollution control game from environmental economics considered in [23]. There, each player represents a country that is, on one hand, seeking to maximing the net revenue from production of goods and services and, on the other hand, minimizing the environmental damage due to pollution. Pollution is assumed to be a proportional by-product of production and therefore the net revenues (gross revenue minus production cost) of each player, e.g., it can be expressed as a function of emissions. The strategy $x_{i}$ of country $i$, which is chosen from the space $\mathcal{X}_{i}=\mathbb{R}$, can be seen as the emissions of country $i$. Each country has its own revenue function $g_{i}:\mathcal{X}_{i}\to\mathbb{R}_{+}$ which is nonnegative and concave and depends only on the emissions of country $i$. However, the environmental damage due to emissions depends on the sum of all combined emissions which is stated by a convex damage cost function $d$. Each country tries to minimize its cost function $f_{i}$ (i.e. to maximize her total welfare given as the difference between the net revenue and the damage cost), which is given by $f_{i}(x):=d(x_{1}+...+x_{N})-g_{i}(x_{i})$. Additionally, two types of constraints can be considered. Firstly, each country $i$ can face an environmental constraint $x_{i}\in[0,T_{i}]$ where $T_{i}\geq 0$ is an exogenously given upper bound on emissions, e.g. from an international treaty. Secondly, countries can agree to jointly fix an upper bound for emissions, which is stated by a polyhedral constraint of the form $\sum_{i=1}^{N}\alpha_{i}x_{i}\leq T$, where $\alpha_{i}\in\mathbb{R}_{+}$ for $i=1,...,N,$ are positive coefficients and $T>0$ is an upper bound. In [23], the assumption of a unique Nash equilibrium is made and only those pollution control games are considered in their study. In contrast, with help of Theorem 3.1 (see also [6, Th. 2.6]) we can characterize the set of Nash equilibria as the set of all Pareto solutions of a certain vector optimization problem, regardless of uniqueness or finiteness of the set of equilibria. Furthermore, with Algorithm 4.8, we are able to numerically approximate the set of Nash equilibria. In the following, we present examples for games with $2$ or $3$ players, and where the set of Nash equilibria is nontrivial.